### v1

# Alternative Methods to SHAP Derived from Properties of Kernels: A Note on Theoretical Analysis

Abstract
This study first derives a general and analytical expression of AFA (Additive Feature Attribution) in terms of the kernel in LIME (Local Interpretable Model-agnostic Explanations). Then, we propose some new AFAs that have appropriate properties of kernels or that coincide with the LS prenucleolus in cooperative game theory. We also revisit existing AFAs such as SHAP (SHapley Additive exPlanations) and re-examine the properties of their kernels.
Keywords SHAP, LIME, Kernel, ES, FESP, LS prenucleolus, XAI, Alternatives to SHAP
1 Introduction
In the field of machine learning, Explainable Artificial Intelligence (XAI) refers to techniques and methods that make the decisions and predictions of machine learning models easier to understand. Among them, AFA (Additive Feature Attribution) is a method that decomposes a model’s prediction into the contributions of individual features. Notably, SHAP (SHapley Additive exPlanations), proposed by [5], which is based on the Shapley value [8] in cooperative game theory, is well-known in this context. Recently, research on SHAP has been rapidly expanding ([4]). To reduce the computational cost of SHAP, various methods such as Tree-SHAP[5] and Fast SHAP [3] have been proposed and applied to actual data (for example, [2]). As an alternative to SHAP, [1] considers ES (Equal Surplus) and FESP (Fair Efficient Symmetric Perturbation), both of which are based on solution concepts in cooperative game theory.
In this study, we investigate the relationship between AFA and the kernel in LIME (Local Interpretable Model-agnostic Explanations) as proposed by [6]. [5] characterizes SHAP in terms of the kernel (Kernel SHAP) and derive the expression of SHAP kernel explicitly. Intriguingly, the properties of the SHAP kernel seem different from those that the LIME kernel is expected to have. More specifically, in LIME, the kernel attaches a large weight as a perturbed sample gets closer to the instance being explained, which is different for that of SHAP. In this note, we first provide a general framework to relate an AFA with its associated kernel by deriving an analytical expression of an AFA in terms of its kernel. Then, we propose some new AFAs that have reasonable properties of kernels or that coincide with the LS prenucleolus in cooperative game theory. We also revisit existing AFAs such as SHAP (SHapley Additive exPlanations) and reexamine the properties of their kernels.
2 Preliminaries
Let t and $n$ be the number of the instances and the number of features, respectively. Suppose $N=\{1,...,n\}$, $T=\{1,...,t\}$. The feature input is a $t\times n$ matrix $X=(X_{1},...X_{j},...,X_{n})$. The $j$th feature vector is $X_{j}=(x_{1,j},...,x_{t,j})^{\prime}$ and, for the $\tau$th instance of interest, the vector of features is $x_{\tau}=(x_{\tau,1},...,x_{\tau,j},...,x_{\tau,n})$. Let $f$ be the original prediction model which takes $x_{\tau}$ and produces a prediction.
For an element of the power set of $N$, which is called a coalition in the cooperative game theory, $S\in 2^{N}$, define $x_{\tau,S}=\{x_{\tau,j}|j\in S\}$. $x_{\tau,S}$ is a vector that consists of features in $S$ at $\tau$th instance. Similarly, for $S\in 2^{N}$, define $X_{S}=\{X_{j}|j\in S\}$.
In cooperative game theory, a characteristic function form game is expressed as $(N,v)$ where $N=\{1,...,n\}$ is the set of players and $v$ is a real-valued function on the power set $2^{N}$. For the $\tau$th instance and any coalition $S\in 2^{N}$, when we define $v_{\tau}:2^{N}\longrightarrow\mathbb{R}$ as in (1), a characteristic function form game $(N,v_{\tau})$ is specified for $\tau$:
$v_{\tau}(S)=E\left[\hat{f}(x_{\tau,S},X_{N\setminus S})\right].$ (1)
$v_{\tau}(S)$ is interpreted as the prediction that $f$ produces for the $\tau$th instance, when (i) features $x_{\tau,j}$ where $j\in S$ are known but (ii) features $x_{\tau,k}$ where $k\in N\setminus S$ are unknown. Note that $v_{\tau}(N)=E\left[\hat{f}(x_{\tau,1},...,x_{\tau,n})\right]=\hat{f}(x_{\tau,1},...,x_{\tau,n})$ and $v_{\tau}(\emptyset)=E\left[\hat{f}(X_{1},...,X_{n})\right]=E\left[\hat{f}(X)\right]$, where the former is the prediction when all features at $\tau$th instances are known and the latter is the prediction when none of the features are known. It should be noted that, while standard cooperative game theory assumes that $v(\emptyset)=0$, this is not necessarily satisfied under this machine learning (ML) setting.
With this setup, Additive Feature attribution (AFA) is the method to decompose $v_{\tau}(N)-v_{\tau}(\emptyset)$ into features at $\tau$, depending on their “contributions.”  More precisely, for a characteristic function form game $(N,v_{\tau})$ associated with the $\tau$th instance and for the feature (player) $j$, define a real-valued function $\Psi_{\tau}(j):N\longrightarrow\mathbb{R}$. We hereafter use $\Psi_{\tau}(j)$ and $\Psi_{\tau,j}$ interchangeably and let $\Psi_{\tau}=(\Psi_{\tau,1},...,\Psi_{\tau,n})$. When $\Psi_{\tau}$ satisfies $\sum_{j\in N}\Psi_{\tau,j}=v_{\tau}(N)-v_{\tau}(\emptyset)$, then $\Psi_{\tau}$ is called Additive Feature Attribution (AFA), denoted by $\Psi_{\tau}^{AFA}$.
3 A Brief Review on LIME and Kernel
Here we review [5] and [6], specifically the parts concerning the relationship between LIME and SHAP. In their notation, $x$ is the original representation of an instance being explained and $z$ is a perturbed sample from $x$. They use a binary vector $x^{\prime}$ and a mapping $x=h_{x}(x^{\prime})$, but in this study, just for simplicity, $x=x^{\prime}$ and $z=z^{\prime}$ i.e., the original instances are simplified ([5]), interpretable ([6]) or binary from the beginning.
[5] considers the following minimization problem (LIME, proposed by [6]).
$\xi(x)=\mathop{\rm arg~{}min}\limits_{g\in G}L(f,g,\pi_{x})+\Omega(g)\ \ \ \ \ where$
  • $f$: the original prediction model.
  • $g$: the explanation model defined as $g(z)=\phi_{0}+\sum_{i=1}^{n}\phi_{i}z_{i}$, where $\phi_{i}\in\mathbb{R}$ and $n$ is the number of the features. Let $G$ be the set of all $g$s and let $\phi=(\phi_{1},...,\phi_{n})\in\mathbb{R}^{n}$.
  • $x$: instance being explained.
  • $z$: perturbed sample from $x$. Let $Z$ be the set of all $z$s, including $x$.
  • $\pi_{x}$: local kernel.
and $L$ is the loss function of the minimization problem and $\Omega(g)$ is a measure of complexity of $g$ (a more complex $g$ is penalized). Note that, regarding the kernel $\pi_{x}$, [6] assumes it increases as the distance between $x$ and $z$ decreases, that is, as $z$ gets closer to $x$, a larger weight is attached to $z$.
Based on this setup, [5] assumes $\Omega(g)=0$ and $L(f,g,\pi_{x})=\sum_{z\in Z}\left[f(z)-g(z)\right]^{2}\pi_{x}(z)$. Therefore, the minimization problem of (3) is:
$\mathop{\rm arg~{}min}\limits_{g\in G}\sum_{z\in Z}\left[f(z)-g(z)\right]^{2}\pi_{x}(z)$ $=$ $\mathop{\rm arg~{}min}\limits_{\phi\in\mathbb{R}^{n}}\sum_{z\in Z}\left[f(z)-\left\{\phi_{0}+\sum_{i=1}^{n}\phi_{i}z_{i}\right\}\right]^{2}\pi_{x}(z)$
$=$ $\mathop{\rm arg~{}min}\limits_{\phi\in\mathbb{R}^{n}}\sum_{z\in Z}\left[\sum_{i=1}^{n}\phi_{i}z_{i}-\left\{f(z)-\phi_{0}\right\}\right]^{2}\pi_{x}(z).$ (2)
Now recall $z$ is perturbed sample from $x$ and $x\in Z$. Therefore, summation over $Z$ in (2) coincides with the summation over $2^{N}$ under our notation, and the summation of $\sum_{i=1}^{n}\phi_{i}z_{i}$ coincides with $\sum_{i\in S}\phi_{i}$. Therefore, under our notation, (2) falls into the following:
$\mathop{\rm arg~{}min}\limits_{\phi\in\mathbb{R}^{n}}\sum_{S\in 2^{N}\setminus\emptyset}\left[\sum_{i\in S}\phi_{i}-\left\{v_{\tau}(S)-v_{\tau}(\emptyset)\right\}\right]^{2}\pi_{x_{\tau}}(S).$ (3)
Furthermore, [5] impose a local accuracy condition (or called efficiency) on this optimization problem: $f(x)=g(x)=\phi_{0}+\sum_{i=1}^{n}\phi_{i}x_{i}$ for $x$. We also impose this condition, so (3) becomes:
$\Psi_{\tau}^{AFA}=\mathop{\rm arg~{}min}\limits_{\phi\in\mathbb{R}^{n}with\sum_{i\in N}\phi_{i}=v_{\tau}(N)-v_{\tau}(\emptyset)}\sum_{S\in 2^{N}\setminus\emptyset}\left[\sum_{i\in S}\phi_{i}-\left\{v_{\tau}(S)-v_{\tau}(\emptyset)\right\}\right]^{2}\pi_{x_{\tau}}(S).$ (4)
For the following analysis, we consider 4 to examine relationship between SHAP and other AFAs alternatives to SHAP and the associated kernel $\pi_{x_{\tau}}(S)$.
4 Results
4.1 Conditions on Kernel
Regarding the kernel $\pi_{x_{\tau}}(S)$ in (4), we impose the following two conditions:
$\sum_{S\subseteq N\setminus\{i,j\}}\pi_{x_{\tau}}(S\cup\{i\})=1\qquad\Big{(}\forall i,j\in N\ with\ i\neq j\Big{)}$ (5)
$\pi_{x_{\tau}}(S)=\pi_{x_{\tau}}(T)\qquad\Big{(}\forall S,T\in 2^{N}\ with\ |S|=|T|\Big{)}$ (6)
(5) is only necessary for normalization and is not essential. (6) states that, in terms of the number of features, when $S$ and $T$ are equidistant from $N$, the kernel must assign the same weight to $S$ and $T$. This can be considered a form of symmetry, which is a naturally acceptable condition.
4.2 Analytical Solution to the Optimization Problem
Here, we derive $\Psi_{\tau}^{AFA}$ in (4) analytically. The Lagrangian of (4) is:
$\mathcal{L}(\phi_{1},...\phi_{n},\lambda)=\sum_{S\in 2^{N}\setminus\{\emptyset\}}\left[\sum_{i\in S}\phi_{i}-\{v_{\tau}(S)-v_{\tau}(\emptyset)\}\right]^{2}\cdot\ \pi_{x_{\tau}}(S)-\lambda\left[\sum_{i\in N}\phi_{i}-v_{\tau}(N)+v_{\tau}(\emptyset)\right].$
The F.O.C. on $\phi_{j}$ is:
$\sum_{S\in 2^{N}\setminus\{\emptyset\}:j\in S}2\left(\sum_{i\in S}\phi_{i}-\{v_{\tau}(S)-v_{\tau}(\emptyset)\}\right)\cdot\ \pi_{x_{\tau}}(S)-\lambda=0.$
Therefore, for any $i,j\in N$ with $i\neq j$, the following holds:
$\sum_{S\in 2^{N}\setminus\{\emptyset\}:i\in S}\left(\sum_{k\in S}\phi_{k}-\{v_{\tau}(S)-v_{\tau}(\emptyset)\}\right)\cdot\pi_{x_{\tau}}(S)=\sum_{S\in 2^{N}\setminus\{\emptyset\}:j\in S}\left(\sum_{k\in S}\phi_{k}-\{v_{\tau}(S)-v_{\tau}(\emptyset)\}\right)\cdot\pi_{x_{\tau}}(S)$
$\Longleftrightarrow$ $\sum_{S\subseteq N\setminus\{i,j\}}\left(\sum_{k\in S\cup\{i\}}\phi_{k}-\{v_{\tau}(S\cup\{i\})-v_{\tau}(\emptyset)\}\right)\cdot\pi_{x_{\tau}}(S\cup\{i\})$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad=\sum_{S\subseteq N\setminus\{i,j\}}\left(\sum_{k\in S\cup\{j\}}\phi_{k}-\{v_{\tau}(S\cup\{j\})-v_{\tau}(\emptyset)\}\right)\cdot\pi_{x_{\tau}}(S\cup\{j\})$
$\Longleftrightarrow$ $\sum_{S\subseteq N\setminus\{i,j\}}\bigg{(}\pi_{x_{\tau}}(S\cup\{i\})\cdot\phi_{i}-\pi_{x_{\tau}}(S\cup\{j\})\cdot\phi_{j}\bigg{)}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad=\sum_{S\subseteq N\setminus\{i,j\}}\bigg{(}\pi_{x_{\tau}}(S\cup\{i\})\cdot\{v_{\tau}(S\cup\{i\})-\pi_{x_{\tau}}(S\cup\{j\})\cdot\{v_{\tau}(S\cup\{j\})\bigg{)}$
$\Longleftrightarrow$ $\phi_{i}-\phi_{j}=\sum_{S\subseteq N\setminus\{i,j\}}\bigg{(}\pi_{x_{\tau}}(S\cup\{i\})\cdot\{v_{\tau}(S\cup\{i\})-\pi_{x_{\tau}}(S\cup\{j\})\cdot\{v_{\tau}(S\cup\{j\})\bigg{)},$
which implies:
$\phi_{1}-\sum_{S:1\in S\neq N}\pi_{x_{\tau}}(S)\cdot v_{\tau}(S)=...=\displaystyle\phi_{n}-\sum_{S:n\in S\neq N}\pi_{x_{\tau}}(S)\cdot v_{\tau}(S).$ (7)
Therefore, $\phi=(\phi_{1},...,\phi_{j},...,\phi_{n})$ that satisfies (7) and $\sum_{j\in N}\phi_{j}=v_{\tau}(N)-v_{\tau}(\emptyset)$ is derived as:
${\phi_{j}=\sum_{S:j\in S\neq N}\pi_{x_{\tau}}(S)\cdot v_{\tau}(S)+\frac{v_{\tau}(N)-v_{\tau}(\emptyset)-\sum_{i\in N}\left\{\sum_{S:i\in S\neq N}\pi_{x_{\tau}}(S)\cdot v_{\tau}(S)\right\}}{n}}.$ (8)
(8) expresses the AFA $\phi$ as a function of the associated kernels $\pi_{x_{\tau}}(S)$ enabling the establishment of an AFA from any kernels. In the following sections, we examine several AFAs, some proposed by existing research, while others are newly proposed and generated by kernels having appropriate properties.
4.3 SHAP
In [5], the kernel of SHAP is specified as follows:
$\pi_{x_{\tau}}(S)=\displaystyle{\frac{n-1}{{}_{n}C_{|S|}\cdot|S|\cdot(n-|S|)}}.$ (9)
Given that the solution of the minimization problem (4) remains unchanged even if a constant in the kernel is altered, we slightly modify the kernel as follows,
$\pi_{x_{\tau}}(S)=\displaystyle{\frac{n}{{}_{n}C_{|S|}\cdot|S|\cdot(n-|S|)}}$ (10)
so that (5) is satisfied. By substituting (10) into (8), we obtain the following:
$\Psi_{\tau,j}^{SHAP}=\phi_{j}=\displaystyle\sum_{S\subseteq N\setminus j}\frac{|S|!(n-|S|-1)!}{n!}\left(v_{\tau}(S\cup\{j\})-v_{\tau}(S)\right).$
That is, SHAP is derived as an AFA generated from the kernel expressed in (10). Therefore, it may be more appropriate to consider (10) rather than (9) as the kernel for SHAP. Additionally, it should be noted that the kernel of (9) or (10) reaches its maximum if $|S|=0$ and $|S|=n$, and it has a concave shape regarding $|S|$, which is different from [6] where the weight assigned by the kernel increases as a perturbed sample gets closer to the instance being explained.
4.4 ES and FESP in [1]
As alternative AFAs to SHAP, [1] proposes ES (Equal Surplus) and FESP (Fair Efficient Symmetric Perturbation), based on the solution concepts in cooperative game theory.
First, consider the following kernel:
$\pi_{x_{\tau}}(S)=\left\{\begin{array}[]{ll}1&\mbox{if}\quad|S|=1\\ 0&\mbox{if}\quad 2\leq|S|\leq n-1\\ \infty&\mbox{if}\quad S=N.\\ \end{array}\right.$ (11)
Similarly to the previous case, by substituting (11) into (8), $\phi_{i}$ becomes as follows:
$\Psi_{\tau,j}^{ES}=\phi_{j}=\displaystyle v_{\tau}(\{j\})+\frac{v_{\tau}(N)-v_{\tau}(\emptyset)-\sum_{k\in N}v_{\tau}(\{k\})}{n}.$
That is, $\phi_{i}$ coincides with ES.
Next, suppose the following kernel:
$\pi_{x_{\tau}}(S)=\left\{\begin{array}[]{ll}w_{\tau}&\mbox{if}\quad|S|=1\\ 0&\mbox{if}\quad 2\leq|S|\leq n-2\\ 1-w_{\tau}&\mbox{if}\quad|S|=n-1\\ \infty&\mbox{if}\quad S=N.\\ \end{array}\right.$ (12)
Then, (8) follows that the associated solution of the minimization problem is FESP:
$\Psi_{\tau,j}^{FESP}=\phi_{j}=\displaystyle w_{\tau}\Big{(}v_{\tau}(\{j\})-v_{\tau}(\emptyset)\Big{)}+(1-w_{\tau})\Big{(}v_{\tau}(\emptyset)-v_{\tau}(N\backslash\{j\})\Big{)}.$
Note that the kernel of (11) and (12) also have a concave shape regarding $|S|$.
4.5 AFA based on LS preucleolus
Consider the following kernel:
$\pi_{x_{\tau}}(S)=\left\{\begin{array}[]{ll}\displaystyle{\frac{1}{2^{n-2}}}&\mbox{if}\quad S\neq N\\ \infty&\mbox{if}\quad S=N.\end{array}\right.$ (13)
Note that the shape of this kernel is not concave with respect to $|S|$, although it is still different from [6] in that the shape is flat unless $S=N$. By substituting (13) into (8), the resulting $\phi_{i}$ is:
$\Psi_{\tau,j}^{PNucl}=\phi_{j}=\displaystyle 2\left(\frac{1}{2^{n-1}}\displaystyle\sum_{S:j\in S}v_{\tau}(S)\right)+\frac{v_{\tau}(N)-v_{\tau}(\emptyset)-\sum_{i\in N}\left\{2\left(\frac{1}{2^{n-1}}\sum_{S:i\in S}v_{\tau}(S)\right)\right\}}{n}$
Intriguingly, this solution is identical to that in the following minimization problem in which a kernel does not appear, thus coinciding with the LS prenucleolus proposed by [7]:
$\mathop{\rm arg~{}min}\limits_{\phi\in R^{n}:\sum_{i\in N}\phi_{i}=v_{\tau}(N)-v_{\tau}(\emptyset)}\sum_{S\in 2^{N}\setminus\emptyset}\left[\sum_{i\in S}\phi_{i}-\{v_{\tau}(S)-v_{\tau}(\emptyset)\}\right]^{2}.$
4.6 AFA with a reasonable kernel (I)
The next kernel we consider is as follows:
$\pi_{x_{\tau}}(S)=\left\{\begin{array}[]{ll}\displaystyle{\frac{|S|}{n\cdot 2^{n-3}}}&\mbox{if}\quad S\neq N\\ \infty&\mbox{if}\quad S=N.\\ \end{array}\right.$ (14)
This kernel satisfies the conditions of (5) and (6). Furthermore, this is increasing in $|S|$ and thus consistent with the condition on the kernel in [6]. By substituting (14) into (8), we have:
$\Psi_{\tau,j}^{LnK}=\displaystyle{\phi_{j}=\sum_{S:j\in S\neq N}\frac{|S|}{n\cdot 2^{n-3}}\cdot v_{\tau}(S)+\frac{v_{\tau}(N)-v_{\tau}(\emptyset)-\sum_{i\in N}\left\{\sum_{S:i\in S\neq N}\frac{|S|}{n\cdot 2^{n-3}}\cdot v_{\tau}(S)\right\}}{n}},$
which is the first AFA we propose as an alternative to SHAP. The superscript $LnK$ stands for linealy increasing kernel.
4.7 AFA with a reasonable kernel (II)
Lastly, assume the following kernel:
$\pi_{x_{\tau}}(S)=\left\{\begin{array}[]{ll}\displaystyle{\frac{2^{|S|-1}}{3^{n-2}}}&\mbox{if}\quad S\neq N\\ \infty&\mbox{if}\quad S=N.\\ \end{array}\right.$
This kernel also satisfies (5) and (6) and is increasing in $|S|$. More specifically, each time $|S|$ increases by $1$, the value of the kernel doubles. Then, we get the following expression, which is our second proposed AFA alternative to SHAP.
$\Psi_{\tau,j}^{ExK}=\displaystyle{\phi_{j}=\sum_{S:j\in S\neq N}\frac{2^{|S|-1}}{3^{n-2}}\cdot v_{\tau}(S)+\frac{v_{\tau}(N)-v_{\tau}(\emptyset)-\sum_{i\in N}\left\{\sum_{S:i\in S\neq N}\frac{2^{|S|-1}}{3^{n-2}}\cdot v_{\tau}(S)\right\}}{n}}$
The superscript $ExK$ stands for exponentially increasing kernel.
5 Conclusion
In this study, we first derive an analytical and general expression of an AFA as a function of its associated kernel. Next, we compute several AFAs based on representations of several different specific kernels. Among the existing AFAs, we show that for SHAP, by slightly modifying the kernel into an appropriate form, the generated AFA coincides with SHAP. Additionally, for ES and FESP, we derive the representations of the corresponding kernels. The last three kernels and the AFAs generated from them are proposed for the first time in this study. $\Psi_{\tau}^{PNucl}$ has a kernel that is not concave and coincides with the notion of the LS prenucleolus in the cooperative game theory. $\Psi_{\tau}^{LnK}$ and $\Psi_{\tau}^{ExK}$ are generated as AFAs from kernels that have desirable properties and consistent with the idea from [6] that the kernel assigns a large weight as a perturbed sample gets closer to the instance being explained.
The extent to which these AFAs show different decomposition patterns in experiments using actual data is an empirical question of great importance and one that should be addressed promptly. Another important theme is how the newly presented $\Psi_{\tau,j}^{LnK}$ and $\Psi_{\tau,j}^{ExK}$ in this study can be characterized from the perspective of cooperative game theory, for example, whether they can be axiomatized, is also worth investigating.